# Building a Balanced k-d Tree in O(kn log n) Time

### v44

@@ 2.3 Results for the $\mathitbf{O\left(kn\;\boldsymbol{\log}\;n\right)}$ Algorithm
The $O\left(kn\log n\right)$ k-d tree-building algorithm was implemented in the Java language, and the single-threaded performance of the merge sorting, duplicate tuple removal, and k-d tree-building methods was measured using a 2.3 GHz Intel i7 processor. Figure 3 shows the total time in seconds that was required to perform the initial merge sorts, remove the duplicate tuples, and build the k-d tree, plotted versus $n\log_{2}\left(n\right)$ for $2^{18}\leq n\leq 2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed line of Figure 3 shows the least-squares fit of the total time $t$ to the function $t=mn\log_{2}\left(n\right)$ where $m$ is the slope of the line. The correlation coefficient $r=0.998$ indicates an adequate least-squares fit; hence, the execution times are proportional to $n\log_{2}\left(n\right)$. The question of whether these execution times are proportional to $k$ will be addressed in Section 4 of this article.
@@
The $O\left(kn\log n\right)$ k-d tree-building algorithm was parallelized via Java threads and its performance was measured for one to eight threads using a 2.3 GHz Intel quad-core i7 processor. Figure 4 shows the total time in seconds that was required to perform the initial merge sorts, remove the duplicate tuples, and build the k-d tree, plotted versus the number of threads $q$ for $n=2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed curve of Figure 4 shows the least-squares fit of the total time $t$ to the equation
@@
This equation will be discussed in Section 4 of this article. The correlation coefficient $r=0.9867$ indicates an acceptable least-squares fit.
@@ 3.1 The $\mathitbf{O\left(n\;\boldsymbol{\log}\;n\right)}$ Algorithm
In order to understand the performance of the $O\left(kn\log n\right)$ k-d tree-building algorithm relative to that of other algorithms, the performance was compared to a $O\left(n\log n\right)$ k-d tree-building algorithm that incorporates a $O\left(n\right)$ median-finding algorithm [6]. Most of the complexity of the $O\left(n\log n\right)$ algorithm is limited to the $O\left(n\right)$ median-finding algorithm. The application of this $O\left(n\log n\right)$ k-d tree-building algorithm to sort $\left(x,y,z\right)$ tuples is described as follows.
@@ 3.2 Results for the $\mathitbf{O\left(n\;\boldsymbol{\log}\;n\right)}$ Algorithm
The $O\left(n\log n\right)$ k-d tree-building algorithm was implemented in the Java language, and the single-threaded performance of the merge sorting, duplicate tuple removal, and k-d tree-building methods was measured using a 2.3 GHz Intel i7 processor. Figure 5 shows the total time in seconds that was required to perform the initial merge sort, remove the duplicate tuples, and build the k-d tree, plotted versus $n\log_{2}\left(n\right)$ for $2^{18}\leq n\leq 2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed line of Figure 5 shows the least-squares fit of the total time $t$ to the function $t=mn\log_{2}\left(n\right)$ where $m$ is the slope of the line. The correlation coefficient $r=0.9986$ indicates an adequate least-squares fit.
@@
The $O\left(n\log n\right)$ k-d tree-building algorithm was parallelized via Java threads and its performance was measured for one to eight threads using a 2.3 GHz Intel quad-core i7 processor. Figure 6 shows the total time in seconds that was required to perform the initial merge sort, remove the duplicate tuples, and build the k-d tree, plotted versus the number of threads $q$ for $n=2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed curve of Figure 6 shows the least-squares fit of the total time $t$ to Equation 1. The correlation coefficient $r=0.9958$ indicates an acceptable least-squares fit.
@@ 4 Discussion
but rather to the model that is expressed by Equation 1 in Section 2.3 of this article
@@
Setting $dt/dq$ to zero in Equation 3 and solving for $q$ predicts that the minimum execution time should occur at $q=\sqrt{t_{1}/m_{\mathrm{c}}}$ threads. Substituting into this equation for $q$ the respective values of $t_{1}$ and $m_{\mathrm{c}}$ that were obtained via least-squares fitting for the $O\left(kn\log n\right)$ and $O\left(n\log n\right)$ algorithms predicts minima at $q=5.87$ and $q=6.15$ threads respectively. These minima are depicted in Figures 4 and 6, which predict decreased performance of both k-d tree building algorithms for greater than eight threads. The decreased performance for as few as eight threads is a consequence of the relatively large values for $m_{\mathrm{c}}$ (1.96 and 1.75 seconds per thread respectively) that were obtained via least-squares fitting. Figure 9 in Section 6.10 plots execution time data obtained using an Intel 6-core i7 processor that supports concurrent execution of up to 12 threads. The least-squares fit to the data demonstrates a minimum execution time at $q=\sqrt{t_{1}/m_{\mathrm{c}}}=\sqrt{5.92/0.0627}=9.72$ threads.
@@
The $O\left(n\right)$ median-finding algorithm is complicated and requires careful implementation to achieve optimum performance. For example, the median-finding algorithm utilizes a sorting algorithm to sort large numbers of five-element arrays. For the initial implementation of the $O\left(n\right)$ median-finding algorithm, these small arrays were sorted via the merge sort algorithm that is used for the initial sort of the index array [15]. That merge sort algorithm is not optimized for sorting small arrays and hence resulted in poor performance of the $O\left(n\right)$ median-finding algorithm and consequent poor performance of the $O\left(n\log n\right)$ k-d tree-building algorithm. Replacing the merge sort algorithm with an insertion sort algorithm that is optimized for sorting small arrays [5] conferred a 30 percent improvement to the performance of the $O\left(n\log n\right)$ k-d tree-building algorithm. (The Appendix describes a better alternative than insertion sort.)

### v45

@@ 2.3 Results for the $\mathitbf{O\left(kn\;\boldsymbol{\log}\;n\right)}$ Algorithm
The $O\left(kn\log n\right)$ k-d tree-building algorithm was implemented in the Java language, and the single-threaded performance of the merge sorting, duplicate tuple removal, and k-d tree-building methods was measured using a 2.3 GHz Intel i7 processor. Figure 3 shows the total time in seconds that was required to perform the initial merge sorts, remove the duplicate tuples, and build the k-d tree, plotted versus $n\log_{2}\left(n\right)$ for $2^{18}\leq n\leq 2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed line depicts the least-squares fit of the total time $t$ to the function $t=mn\log_{2}\left(n\right)$ where $m$ is the slope of the line. The correlation coefficient $r=0.998$ indicates an adequate least-squares fit; hence, the execution times are proportional to $n\log_{2}\left(n\right)$. The question of whether these execution times are proportional to $k$ is discussed in Section 4 of this article.
@@
The $O\left(kn\log n\right)$ k-d tree-building algorithm was parallelized via Java threads and its performance was measured for one to eight threads using a 2.3 GHz Intel quad-core i7 processor. Figure 4 shows the total time in seconds that was required to perform the initial merge sorts, remove the duplicate tuples, and build the k-d tree, plotted versus the number of threads $q$ for $n=2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed curve depicts the least-squares fit of the total time $t$ to the equation
@@
The correlation coefficient $r=0.9867$ indicates an acceptable least-squares fit. Equation 1 and the least-squares fit are discussed in Sections 4 and 6.10 respectively.
@@ 3.1 The $\mathitbf{O\left(n\;\boldsymbol{\log}\;n\right)}$ Algorithm
In order to understand the performance of the $O\left(kn\log n\right)$ k-d tree-building algorithm relative to that of other algorithms, the performance was compared to a $O\left(n\log n\right)$ k-d tree-building algorithm that incorporates a $O\left(n\right)$ median-finding algorithm [6]. The application of the $O\left(n\log n\right)$ k-d tree-building algorithm to sort $\left(x,y,z\right)$ tuples is described as follows.
@@ 3.2 Results for the $\mathitbf{O\left(n\;\boldsymbol{\log}\;n\right)}$ Algorithm
The $O\left(n\log n\right)$ k-d tree-building algorithm was implemented in the Java language, and the single-threaded performance of the merge sorting, duplicate tuple removal, and k-d tree-building methods was measured using a 2.3 GHz Intel i7 processor. Figure 5 shows the total time in seconds that was required to perform the initial merge sort, remove the duplicate tuples, and build the k-d tree, plotted versus $n\log_{2}\left(n\right)$ for $2^{18}\leq n\leq 2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed line depicts the least-squares fit of the total time $t$ to the function $t=mn\log_{2}\left(n\right)$ where $m$ is the slope of the line. The correlation coefficient $r=0.9986$ indicates an adequate least-squares fit.
@@
The $O\left(n\log n\right)$ k-d tree-building algorithm was parallelized via Java threads and its performance was measured for one to eight threads using a 2.3 GHz Intel quad-core i7 processor. Figure 6 shows the total time in seconds that was required to perform the initial merge sort, remove the duplicate tuples, and build the k-d tree, plotted versus the number of threads $q$ for $n=2^{24}$ $\left(x,y,z,w\right)$ tuples of randomly-generated 32-bit integers. The dashed curve depicts the least-squares fit of the total time $t$ to Equation 1. The correlation coefficient $r=0.9958$ indicates an acceptable least-squares fit.
@@ 4 Discussion
but rather to the model that is expressed by Equation 1
@@
Setting $dt/dq$ to zero in Equation 3 and solving for $q$ predicts that the minimum execution time should occur at $q=\sqrt{t_{1}/m_{\mathrm{c}}}$ threads. Substituting into this equation for $q$ the respective values of $t_{1}$ and $m_{\mathrm{c}}$ that were obtained via least-squares fitting for the $O\left(kn\log n\right)$ and $O\left(n\log n\right)$ algorithms predicts minima at $q=5.87$ and $q=6.15$ threads respectively. These minima are depicted in Figures 4 and 6 that predict decreased performance of both k-d tree building algorithms for greater than eight threads. The decreased performance for as few as eight threads is a consequence of the relatively large values for $m_{\mathrm{c}}$ (1.96 and 1.75 seconds per thread respectively) that were obtained via least-squares fitting. Figure 9 in Section 6.10 plots execution time data obtained using an Intel 6-core i7 processor that supports concurrent execution of two threads per core. The least-squares fit to the data demonstrates a minimum execution time at $q=\sqrt{t_{1}/m_{\mathrm{c}}}=\sqrt{5.92/0.0627}=9.72$ threads.
@@
The $O\left(n\right)$ median-finding algorithm is complicated and requires careful implementation to achieve optimum performance. For example, the median-finding algorithm utilizes a sorting algorithm to sort large numbers of five-element arrays. For the initial implementation of the $O\left(n\right)$ median-finding algorithm, these small arrays were sorted via the merge sort algorithm that is used for the initial sort of the index array [15]. That merge sort algorithm is not optimized for sorting small arrays and hence resulted in poor performance of the $O\left(n\right)$ median-finding algorithm and consequent poor performance of the $O\left(n\log n\right)$ k-d tree-building algorithm. Replacing the merge sort algorithm with an insertion sort algorithm that is optimized for sorting small arrays [5] conferred a 30 percent improvement to the performance of the $O\left(n\log n\right)$ k-d tree-building algorithm. Section 6.5 describes a preferred alternative to insertion sort.